OpenStudy (anonymous):

If 2/3 = x/y, then which of the following must be true? a. (2 + x)/3 = (x + 2)/y b. 3/2 = x/y c. (2 + 3)/3 = (x + y)/y d. (2 + 1)/3 = (x + 1)/y

OpenStudy (anonymous):

c is true. \(\frac{2}{3} = \frac{x}{y}\) so \(\frac{2y}{3} = x\) if we multiply across by y. Sub this value in for x in equation 3 to show it.

OpenStudy (anonymous):

Thank you!